A metallurgist has one alloy containing 21% copper and another containing 60% copper. How many pounds of each alloy must he use to make 54 pounds of a third alloy containing 27% copper?
Answer by richwmiller(17219)

a+b=54,
21*a+60*b=27*54
a=54-b
21*(54-b)+60*b=1458
1134-21b+60*b=1458
39*b=324
b=8.30769231
a=54-b
a=45.6923077 lb. at 21% b=8.30769231 lb. at 60%
check
21*45.6923077+60*8.30769231=27*54
959.538462+498.461538=1458
1458.0=1458
ok